OpenStudy (anonymous):

Choose the correct simplification of the expression (2ab5)3? help please

OpenStudy (harsimran_hs4):

is the expression (2a(b^5))^3 ?

OpenStudy (anonymous):

(2ab^5)^3

OpenStudy (harsimran_hs4):

ok.. property : (abc)^d = (a^d)(b^d)(c^d) apply this property and tell me what you get ?

OpenStudy (anonymous):

i have no idea

OpenStudy (harsimran_hs4):

ok i will do the first step for you (2ab^5)^3 = (2^3) * (a^3) * ((b^5))^3 where * = multiplication clear ?

OpenStudy (anonymous):

good thanks :)